3. What is the scale elasticity of F(K,L) K22L, where K denotes capital and L denotes labor? What relationship between capital and labor must hold for scale elasticity to equal one?
Answer #1

Q = F(K, L) = K2 + 2L0.5

(1)

Scale elasticity (E) = (\partialQ/\partialK) x (K/Q) + (\partialQ/\partialL) x (L/Q)

= 2K x (K/Q) + [(2 x 0.5) / L0.5] x (L/Q)

= (2K2/Q) + (L0.5/Q)

= (1/Q) x (2K2 + L0.5)

= (2K2 + L0.5) / (K2 + 2L0.5)

(2)

When E = 1,

(2K2 + L0.5) = (K2 + 2L0.5)

K2 = L0.5

Squaring,

K4 = L (Required condition)